An end-of-aisle price promotion changes the price elasticity of a good from −2 to −3. Suppose the normal price is $40, which equates marginal revenue with marginal cost at the initial elasticity of –2.

What should the promotional price be when the elasticity changes to –3? (Hint: In other words, what price will equate marginal revenue and marginal cost?)

A. $39.00

B. $30.00

C. $42.00

D. $27.00
Answer #1

Firstly calculate MC when e = -2, where MR = MC

(P-MC) / P = 1 / IeI

Here P = $40 and e = -2

(40 - MC) / 40 = 1/ I-2I

(40 - MC) / 40 = 1 / 2

80 - 2MC = 40

40 = 2MC

MC = 40

Now, as we have MC, we will calculate the new price when e = -3

(P-MC) / P = 1 / IeI

(P - 20) / P = 1 / I-3I

(P - 20) / P = 1 / 3

3P - 60 = P

2P = 60

P = 30